OpenStudy (sissyedgar):

The total snowfall in Burke, Virginia, this January was 7 1/4 in. In February, 20% more snow fell than in January. How much snow fell in February in Burke, Virginia? Express your answer as a mixed number in simplest form. ( ) ( ) ______________in. ( )

OpenStudy (sissyedgar):

??? I need serious help in serious time!

OpenStudy (sissyedgar):

@MonsterMonkey

OpenStudy (snowsurf):

You know January there was 7 1/4 and there was 20% more snow in February. First turn the fraction into a improper fraction so 7 1/4 will become 29/4\[\frac{ 29 }{ 4 } + \frac{ 29 }{ 4 }(\frac{ 20 }{ 100 })\] 20/100 can be reduced to 1/5 so it becomes \[\frac{ 29 }{ 4 } + \frac{ 29 }{ 4 }(\frac{ 1 }{ 5 })\] \[\frac{ 29 }{ 4 } + \frac{ 29 }{ 20 } = \frac{ 29 }{ 20 } + \frac{ 145 }{ 20 } = \frac{ 174 }{ 20 }\] When you reduce the fraction you should get \[8\frac{ 7 }{ 10 }\] for the final answer.

OpenStudy (sissyedgar):

ty